The measure of two adjacent angles forms an angle of 143. The measure of the smaller angle is seven degrees more than one-third the measure of the larger angle. What are the measures of the two angles?

Answer by KnightOwlTutor(293)

x=larger angle
1/3x+7=smaller angle
x+1/3x+7=143
Subtract 7 from both sides

x+x1/3=136
1+1/3=4/3
4/3x=136

multiply each side by 3/4

x=(136)(3/4)
x=102
Let's doublecheck our answer
smaller angle is (1/3)(102)+7=41

102+41=143